# Automorphisms of relatively hyperbolic groups and the Farrell–Jones Conjecture

### 23 Nov 2023 v1

@@ 5.B. The Periodic JSJ tree
Let $G$ be a one-ended hyperbolic group relative to ${\mathcal{P}}$ and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$. If for every $[P]\in{\mathcal{P}}$ the group $P_{\Phi}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G_{\Phi}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
Consider the action of $G_{\Phi}$ on $T^{\mathrm{Per}}$. By Corollary 5.20, the action of $G_{\Phi}$ on $T^{\mathrm{Per}}$ is acylindrical. Up to taking a power of $\Phi$ (which does not change the conclusion by 2.1), we may assume that $\phi=[\Phi]$ is almost rotationless.
@@
Suppose first that $G_{v}$ is elementary. If $G_{v}\subseteq P$ for some $[P]\in{\mathcal{P}}$, then $(G_{\Phi})_{v}$ is a subgroup of $P_{\Phi}$. In particular, it belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by Theorem 2.1 $(1)$. If $G_{v}$ is infinite virtually cyclic, then $(G_{\Phi})_{v}$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
$1\to F\to(G_{\Phi})_{v}\to\pi_{1}(\Sigma_{v})\rtimes\mathbb{Z}\to 1$
The groups $F$ and $\pi_{1}(\Sigma_{v})\rtimes\mathbb{Z}$ belong to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. For the latter one may use that surface bundles over the circle are locally $\mathrm{CAT}(0)$ and apply [Weg12]. Moreover, for every virtually cyclic group $Q\subseteq\pi_{1}(\Sigma_{v})\rtimes\mathbb{Z}$, the preimage of $Q$ in $(G_{\Phi})_{v}$ is virtually cyclic, and hence belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. Thus, the group $(G_{\Phi})_{v}$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
Suppose that $G_{v}$ is rigid. By 5.19, since $\phi$ is almost rotationless, there exists $[H]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ such that $G_{v}=H$.
Suppose that $\Phi\in\phi$ is such that $\mathrm{Per}(\Phi)=H$. Then $(G_{\Phi})_{v}$ is isomorphic to $\mathrm{Per}(\Phi)\rtimes_{\Phi}\mathbb{Z}$. By Lemma 2.6, we have $(G_{\Phi})_{v}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@ 5.C. An aside on slender peripherals
We also isolate here an interesting consequence of 5.19 for automorphisms of groups hyperbolic relative to slender groups. Recall that a group is slender if all its subgroups are finitely generated.
@@
See also the proof of [GL15, Theorem 8.2]. We claim that it suffices to prove that $\mathrm{Per}(\Phi)$ is finitely generated. Indeed, suppose that $\mathrm{Per}(\Phi)$ is generated by $a_{1},\ldots,a_{n}$. For every $i\in\{1,\ldots,n\}$, let $k_{i}$ be such that $\Phi^{k_{i}}(a_{i})=a_{i}$. Let $N=k_{1}\ldots k_{n}$. Then $\mathrm{Per}(\Phi)=\operatorname{\mathrm{Fix}}(\Phi^{N})$.
@@
Let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G)$. Suppose first that $G$ is one-ended relative to ${\mathcal{H}}=\mathrm{Per}_{\mathrm{NP}}(\phi)$. Let $T^{\mathrm{Per}}$ be the associated JSJ tree.
@@
By 5.19, the group $\mathrm{Per}(\Phi)$ is equal to the stabiliser of a vertex of $T^{\mathrm{Per}}$, hence is finitely generated.
Suppose now that $G$ is not one-ended relative to ${\mathcal{H}}$ and consider a minimal reduced Stallings-Dunwoody decomposition $S$ of $G$ such that, for every $[H]\in{\mathcal{P}}\cup{\mathcal{H}}$, the group $H$ is contained in the stabiliser of a vertex of $S$. Recall that edge stabilisers in $S$ are all finite. Since $\mathrm{Per}(\Phi)$ is infinite, it fixes a unique vertex $v$. Since the deformation space of $S$ (see [GL07]) is invariant by $\Phi$, the group $\Phi(G_{v})$ also fixes a unique vertex $w$ in $S$. As $\mathrm{Per}(\Phi)\subseteq\Phi(G_{v})$ and as $\mathrm{Per}(\Phi)$ only fixes $v$, we see that $v=w$. This shows that $\Phi(G_{v})=G_{v}$.
Note that, by minimality of $S$, the group $G_{v}$ is one-ended hyperbolic relative to the restriction ${\mathcal{P}}_{v}$ of ${\mathcal{P}}$ in $G_{v}$. Since ${\mathcal{P}}$ is a set of conjugacy classes of slender groups, so is ${\mathcal{P}}_{v}$. Thus, the conclusion follows from the one-ended case applied to the restriction $\Phi|_{G_{v}}$. ∎
We remark that Minasyan–Osin [MO12, Corollary 1.3] also proved that the fixed subgroup of the automorphism of any hyperbolic group relative to slender groups is finitely generated.
@@ 6. A combination theorem for the Farrell–Jones conjecture
Let $G$ be a finitely generated group and let $\Phi\in\operatorname{\mathrm{Aut}}(G)$. If $F$ is a malnormal subgroup of $G$ whose conjugacy class is $\Phi$-periodic, recall the definition of $F_{\Phi}$ from just above Corollary 5.21. In this section, we prove the following combination theorem.
@@
Let $G=G_{1}\ast\ldots\ast G_{k}\ast F_{N}$ be a free product of finitely generated groups, let =${\mathcal{F}}^{\prime}=\{[G_{1}],\ldots,[G_{k}]\}$ and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}}^{\prime})$. If for each $i\in\{1,\ldots,k\}$, the group $(G_{i})_{\Phi}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G\rtimes_{\Phi}\mathbb{Z}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@ 6.A. The nonsporadic case
Let $G=G_{1}\ast\ldots\ast G_{k}\ast F_{N}$ be a free product of groups, let =${\mathcal{F}}=\{[G_{1}],\ldots,[G_{k}]\}$ and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ be fully irreducible. If for each $i\in\{1,\ldots,k\}$, the group $(G_{i})_{\Phi}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G\rtimes_{\Phi}\mathbb{Z}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
Let $S$ be the Grushko $(G,{\mathcal{F}})$-tree given by Lemma 3.3. Let ${\mathcal{P}}_{S}(\Phi)$ be the .$\lVert.\rVert_{S}$-maximal polynomial subgroups of $\Phi$. By Theorem 3.2, up to taking a power of $\Phi$ (which does not change the conclusion of Lemma 6.1 by Theorem 2.1 $(2)$) the group $G\rtimes_{\Phi}\mathbb{Z}$ is hyperbolic relative to the suspension of ${\mathcal{P}}_{S}(\Phi)$. By Proposition 3.4, for every $[P]\in{\mathcal{P}}_{S}(\Phi)$, either $[P]\in{\mathcal{F}}$ or $P$ is infinite cyclic. In either case $P_{\Phi}$ is contained in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. By Theorem 2.4, the group $G\rtimes_{\Phi}\mathbb{Z}$ is contained in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. ∎
@@ 6.B. The sporadic case
This section follows [BFW23, Proof of Proposition 4.1]. Let $(G,{\mathcal{F}})$ be a sporadic free product and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$. Since $(G,{\mathcal{F}})$ is sporadic, the automorphism $\Phi$ induces a $G$-equivariant homeomorphism of the Bass-Serre tree $T_{\mathcal{F}}$ associated with ${\mathcal{F}}$. This induces an action of $G\rtimes_{\Phi}\mathbb{Z}$ on $T_{\mathcal{F}}$. However, this action is not necessarily acylindrical. In order to apply Theorem 2.5, we will consider the action of $G\rtimes_{\Phi}\mathbb{Z}$ on the tree of cylinders of $T_{\mathcal{F}}$ associated with an admissible relation that we now describe.
@@
Let $Y$ be the cylinder of $T_{\mathcal{F}}$ containing $e$. The stabiliser of $Y$ in $G\rtimes_{\Phi}\mathbb{Z}$ is isomorphic to $\langle\operatorname{\mathrm{Fix}}(\Phi^{n})\rangle_{n\in\mathbb{N}}\rtimes_{\Phi}\mathbb{Z}$.
@@
The proof is identical to [BFW23, Lemma 4.6]. Let $v,v^{\prime}\in VT_{c}$ with $d_{T_{c}}(v,v^{\prime})\geq 6$. We may suppose that $v$ and $v^{\prime}$ correspond to vertices $w$ and $w^{\prime}$ in $T_{\mathcal{F}}$ up to considering adjacent vertices in the path between them. Even after this operation, we have $d_{T_{c}}(v,v^{\prime})\geq 4$.
@@
Let $(G,{\mathcal{F}})$ be a sporadic free product of groups and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$. If for each $[A]\in{\mathcal{F}}$, the group $A_{\Phi}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G\rtimes_{\Phi}\mathbb{Z}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
Let $T_{\mathcal{F}}$ be the Bass-Serre tree associated with ${\mathcal{F}}$ and let $T_{c}$ be its tree of cylinders relative to the commensurability relation. We want to apply Theorem 2.5 to the action of $G\rtimes_{\Phi}\mathbb{Z}$ on $T_{c}$. This action is acylindrical by Lemma 6.3. Thus, it suffices to prove that every vertex stabiliser belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
Suppose now that $v\in V_{1}T_{c}$. By Lemma 6.2, the stabiliser of $v$ is isomorphic to $\mathrm{Per}(\Phi)\rtimes_{\Phi}\mathbb{Z}$. By Lemma 2.6, the group $\mathrm{Per}(\Phi)\rtimes_{\Phi}\mathbb{Z}$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.

### 02 Mar 2026 v2

@@ 5.B. The Periodic JSJ tree
Let $G$ be a hyperbolic group relative to $\mathcal{P}$}. Let ${\mathcal{H}}_{0}$ be a finite set of conjugacy classes of finitely generated subgroups of $G$ such that G is one-ended relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}$. Let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}_{0}^{(t)})$.
If for every $P]\in{\mathcal{P}}$ the group $P_{\Phi}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G_{\Phi}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
Consider the action of $G_{\Phi}$ on $T^{\mathrm{Per}}$. By Corollary 5.22, the action of $G_{\Phi}$ on $T^{\mathrm{Per}}$ is acylindrical. Up to taking a power of $\Phi$ (which does not change the conclusion by ˜2.1), we may assume that $\phi=[\Phi]$ is almost rotationless.
@@
Suppose first that $G_{v}$ is elementary. If $G_{v}\subseteq P$ for some $P]\in{\mathcal{P}}$, then $(G_{\Phi})_{v}$ is a subgroup of $P_{\Phi}$. In particular, it belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by Theorem 2.1 $(1)$. If $G_{v}$ is infinite virtually cyclic, then $(G_{\Phi})_{v}$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
$1\to F\to(G_{\Phi})_{v}\to\pi_{1}(\Sigma_{v})\rtimes\mathbb{Z}\to 1.$
The groups $F$ and $\pi_{1}(\Sigma_{v})\rtimes\mathbb{Z}$ belong to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. Indeed, the Farrell–Jones Conjecture holds for all fundamental groups of (not necessarily compact) connected manifolds (possibly with boundary) of dimension at most $3$, by [41, Theorem 16.1(i)(e)]. As explained in loc. cit. the result is deduced from Corollary 1.3 and Remark 1.4 of [1] where [47, 46] is used. Moreover, for every virtually cyclic group $Q\subseteq\pi_{1}(\Sigma_{v})\rtimes\mathbb{Z}$, the preimage of $Q$ in $(G_{\Phi})_{v}$ is virtually cyclic, and hence belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. Thus, the group $(G_{\Phi})_{v}$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
Suppose that $G_{v}$ is rigid. By ˜5.21, since $\phi$ is almost rotationless, there exists $H]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ such that $G_{v}=H$.
Suppose that $\Phi\in\phi$ is such that $\mathrm{Per}(\Phi)=H$. Then $(G_{\Phi})_{v}$ is isomorphic to $\mathrm{Per}(\Phi)\rtimes_{\Phi}\mathbb{Z}$. By Lemma˜2.6, we have $(G_{\Phi})_{v}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@ 5.C. An aside on slender peripherals
We also isolate here an interesting consequence of ˜5.21 for automorphisms of groups hyperbolic relative to slender groups. Recall that a group is slender if all its subgroups are finitely generated.
@@
See also the proof of [27, Theorem 8.2]. We claim that it suffices to prove that $\mathrm{Per}(\Phi)$ is finitely generated. Indeed, suppose that $\mathrm{Per}(\Phi)$ is generated by $a_{1},\ldots,a_{n}$. For every $i\in\{1,\ldots,n\}$, let $k_{i}$ be such that $\Phi^{k_{i}}(a_{i})=a_{i}$. Let $N=k_{1}\ldots k_{n}$. Then $\mathrm{Per}(\Phi)=\operatorname{\mathrm{Fix}}(\Phi^{N})$.
@@
Let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G)$ and let ${\mathcal{H}}=\mathrm{Per}_{\mathrm{NP}}(\phi)$. Let $({\mathcal{H}}_{i})_{i\geq 1}$ be an increasing sequence of finite set of conjugacy classes of finitely generated subgroups of $G$ such that ${\mathcal{H}}=\bigcup_{i\geq 1}{\mathcal{H}}_{i}$. By [45], for every $i\geq 1$, the group $G$ is finitely presented relative to ${\mathcal{P}}\cup{\mathcal{H}}_{i}$. Thus, the Stallings–Dunwoody JSJ deformation space $D_{i}$ of $G$ over finite groups relative to ${\mathcal{P}}\cup{\mathcal{H}}_{i}$ exists (see [25]).
We claim that there exists $i_{0}$ such that, for every $i\geq i_{0}$, we have $D_{i}=D_{i_{0}}$. The proof is identical to [30, Lemma 8.2] replacing Linnell’s accessibility Theorem by [25, Proposition 4.2]. We sketch the argument. Let $i\geq 1$. By [30, Lemma 2.15], since edge groups are all finite and hence all universally elliptic, for every $T_{i}\in D_{i}$, there exists a sequence of refinements of trees $T_{i}^{1}\to\ldots\to T_{i}^{i-1}\to T_{i}^{i}=T_{i}$ such that, for every $j\leq i$, we have $T_{i}^{j}\in D_{j}$. Let $j\leq i$. Notice that, since edge groups are finite, $T_{i}^{j-1}$ and $T_{i}^{j}$ are not in the same deformation space if and only if they do not have the same infinite elliptic subgroups if and only if the number of orbits of vertices with infinite stabiliser is greater in $T_{i}^{j-1}$ than in $T_{i}^{j}$. By [25, Proposition 4.2], the number of orbits of vertices of $T_{i}^{1}$ with infinite stabiliser is bounded by a constant $J$ depending only on $D_{1}$. Thus, the number of $j\leq i$ such that $T_{i}^{j-1}$ and $T_{i}^{j}$ are not in the same deformation space is bounded by $J$. This forces the spaces $D_{i}$ to eventually stabilise, hence the claim. Let $D_{i_{0}}$ be the stabilised deformation space. As ${\mathcal{P}}\cup{\mathcal{H}}={\mathcal{P}}\cup\bigcup_{i\geq 1}{\mathcal{H}}_{i}$, this shows that any tree in $D_{i_{0}}$ is a $({\mathcal{P}}\cup{\mathcal{H}})$-tree with finite edge groups.
Suppose first that $G$ is one-ended relative to $\mathcal{P}$$\cup{\mathcal{H}}$. Then $D_{i_{0}}$ consists of the trivial tree and $G$ also one-ended relative to ${\mathcal{P}}\cup{\mathcal{H}}_{i_{0}}$. Moreover, since ${\mathcal{H}}_{i_{0}}$ is a finite set and all groups whose conjugacy class is in ${\mathcal{H}}_{i_{0}}$ are finitely generated, up to taking a power of $\phi$, we have $\phi\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}_{i_{0}}^{(t)})$. Let $T^{\mathrm{Per}}$ be the JSJ tree associated with ${\mathcal{P}}\cup{\mathcal{H}}_{i_{0}}\cup{\mathcal{H}}$.
@@
By ˜5.21, the group $\mathrm{Per}(\Phi)$ is equal to the stabiliser of a vertex of $T^{\mathrm{Per}}$, hence is finitely generated.
Suppose now that $G$ is not one-ended relative to $\mathcal{P}$$\cup{\mathcal{H}}$ and consider a tree $S$ in $D_{i_{0}}$. Recall that edge stabilisers in $S$ are all finite. Since $\mathrm{Per}(\Phi)$ is infinite, it fixes a unique vertex $v$. Since the deformation space of $S$ is invariant by $\Phi$, the group $\Phi(G_{v})$ also fixes a unique vertex $w$ in $S$. As $\mathrm{Per}(\Phi)\subseteq\Phi(G_{v})$ and as $\mathrm{Per}(\Phi)$ only fixes $v$, we see that $v=w$. This shows that $\Phi(G_{v})=G_{v}$.
Note that, by minimality of $S$, the group $G_{v}$ is hyperbolic relative to the restriction ${\mathcal{P}}_{v}$ of $\mathcal{P}$} in $G_{v}$ and one-ended relative to ${\mathcal{P}}_{v}\cup({\mathcal{H}}_{i_{0}})_{v}\cup{\mathcal{H}}_{v}$. Since $\mathcal{P}$} is a set of conjugacy classes of slender groups, so is ${\mathcal{P}}_{v}$. Thus, the conclusion follows from the one-ended case applied to the restriction $\Phi|_{G_{v}}$. ∎
We remark that Minasyan–Osin [44, Corollary 1.3] also proved that the fixed subgroup of the automorphism of any hyperbolic group relative to slender groups is finitely generated.
@@ 6. A combination theorem for the Farrell–Jones conjecture
Let $G$ be a finitely generated group and let $\Phi\in\operatorname{\mathrm{Aut}}(G)$. If $F$ is a malnormal subgroup of $G$ whose conjugacy class is $\Phi$-periodic, recall the definition of $F_{\Phi}$ from just above Corollary˜5.23. In this section, we prove the following combination theorem.
@@
Let $G=G_{1}\ast\ldots\ast G_{k}\ast F_{N}$ be a free product of finitely generated groups, let ${\mathcal{F}}^{\prime}=\{[G_{1}],\ldots,[G_{k}]\}$ and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}}^{\prime})$. If for each $i\in\{1,\ldots,k\}$, the group $(G_{i})_{\Phi}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G\rtimes_{\Phi}\mathbb{Z}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@ 6.A. The nonsporadic case
Let $G=G_{1}\ast\ldots\ast G_{k}\ast F_{N}$ be a free product of groups, let ${\mathcal{F}}=\{[G_{1}],\ldots,[G_{k}]\}$ and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ be fully irreducible. If for each $i\in\{1,\ldots,k\}$, the group $(G_{i})_{\Phi}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G\rtimes_{\Phi}\mathbb{Z}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
Let $S$ be the Grushko $(G,{\mathcal{F}})$-tree given by Lemma 3.3. Let ${\mathcal{P}}_{S}(\Phi)$ be the $\lVert.\rVert_{S}$-maximal polynomial subgroups of $\Phi$. By Theorem 3.2, up to taking a power of $\Phi$ (which does not change the conclusion of Lemma 6.1 by Theorem 2.1 $(2)$) the group $G\rtimes_{\Phi}\mathbb{Z}$ is hyperbolic relative to the suspension of ${\mathcal{P}}_{S}(\Phi)$. By Proposition 3.4, for every $P]\in{\mathcal{P}}_{S}(\Phi)$, either $P]\in{\mathcal{F}}$ or $P$ is infinite cyclic. In either case $P_{\Phi}$ is contained in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. By Theorem 2.4, the group $G\rtimes_{\Phi}\mathbb{Z}$ is contained in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. ∎
@@ 6.B. The sporadic case
This section follows [11, Proof of Proposition 4.1]. Let $(G,{\mathcal{F}})$ be a sporadic free product and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$. Since $(G,{\mathcal{F}})$ is sporadic, the automorphism $\Phi$ induces a $G$-equivariant homeomorphism of the Bass–Serre tree $T_{\mathcal{F}}$ associated with $\mathcal{F}$}. This induces an action of $G\rtimes_{\Phi}\mathbb{Z}$ on $T_{\mathcal{F}}$. However, this action is not necessarily acylindrical. In order to apply Theorem 2.5, we will consider the action of $G\rtimes_{\Phi}\mathbb{Z}$ on the tree of cylinders of $T_{\mathcal{F}}$ associated with an admissible relation that we now describe.
@@
Let $Y$ be the cylinder of $T_{\mathcal{F}}$ containing $e$. The stabiliser of $Y$ in $G\rtimes_{\Phi}\mathbb{Z}$ is isomorphic to $\langle\operatorname{\mathrm{Fix}}(\Phi^{n})\rangle_{n\geq 1}\rtimes_{\Phi}\mathbb{Z}$.
@@
The proof is identical to [11, Lemma 4.6]. Let $v,v^{\prime}\in VT_{c}$ with $d_{T_{c}}(v,v^{\prime})\geq 6$. We may suppose that $v$ and $v^{\prime}$ correspond to vertices $w$ and $w^{\prime}$ in $T_{\mathcal{F}}$ up to considering adjacent vertices in the path between them. Even after this operation, we have $d_{T_{c}}(v,v^{\prime})\geq 4$.
@@
Let $(G,{\mathcal{F}})$ be a sporadic free product of groups and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$. If for each $A]\in{\mathcal{F}}$, the group $A_{\Phi}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G\rtimes_{\Phi}\mathbb{Z}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
Let $T_{\mathcal{F}}$ be the Bass–Serre tree associated with $\mathcal{F}$} and let $T_{c}$ be its tree of cylinders relative to the commensurability relation. We want to apply Theorem 2.5 to the action of $G\rtimes_{\Phi}\mathbb{Z}$ on $T_{c}$. This action is acylindrical by Lemma 6.3. Thus, it suffices to prove that every vertex stabiliser belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
Suppose now that $v\in V_{1}T_{c}$. By Lemma 6.2, the stabiliser of $v$ is isomorphic to $\mathrm{Per}(\Phi)\rtimes_{\Phi}\mathbb{Z}$. By Lemma˜2.6, the group $\mathrm{Per}(\Phi)\rtimes_{\Phi}\mathbb{Z}$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.